# Positive Semidefinite Supermartingales and Randomized Matrix Concentration Inequalities

### v3

@@ 4.1 Matrix-Valued Supermartingales
1. For any $\mathbf{v}\in\mathbb{R}^{d}$, the scalar process $\{\mathbf{v}^{\mathsf{T}}\mathbf{Y}_{n}\mathbf{v}\}$ is a supermartingale on $\{\mathcal{F}_{n}\}$ and therefore so is the stopped process $\{\mathbf{v}^{\mathsf{T}}\mathbf{Y}_{n\wedge\tau}\mathbf{v}\}$ due to the optional stopping theorem (e.g. Theorem 10.15 of Klenke [2013]). This implies that $\{\mathbf{Y}_{n\wedge\tau}\}$ is a matrix supermartingale on $\{\mathcal{F}_{n}\}$. 2. For any $\mathbf{v}\in\mathbb{R}^{d}$, we apply the optional sampling theorem (e.g. Theorem 10.11(i) of Klenke [2013]) to the scalar-valued supermartingale $\{\mathbf{v}^{\mathsf{T}}\mathbf{Y}_{n}\mathbf{v}\}$ to obtain that $\mathbf{v}^{\mathsf{T}}(\mathbb{E}\mathbf{Y}_{\tau})\mathbf{v}=\mathbb{E}(\mathbf{v}^{\mathsf{T}}\mathbf{Y}_{\tau}\mathbf{v})\leqslant\mathbb{E}(\mathbf{v}^{\mathsf{T}}\mathbf{Y}_{0}\mathbf{v})=\mathbf{v}^{\mathsf{T}}(\mathbb{E}\mathbf{Y}_{0})\mathbf{v}$. Therefore $\mathbb{E}\mathbf{Y}_{\tau}\preceq\mathbb{E}\mathbf{Y}_{0}$. 3. Apply Theorem 4.2 to the $\mathcal{S}_{d}^{+}$-valued stopped supermartingale $\{\mathbf{Y}_{\tau\wedge n}\}$, we have $\mathbb{E}\mathbf{Y}_{\tau\wedge\infty}\preceq\mathbb{E}\mathbf{Y}_{\tau\wedge 0}$, i.e., $\mathbb{E}\mathbf{Y}_{\tau}\preceq\mathbb{E}\mathbf{Y}_{0}$. ∎
This gives rise to the matrix versions of (VI) and (UVI) via some manipulation of crossing events.
Theorem 4.4 ((Uniformly Randomized) Matrix Ville’s Inequality).
Let $\{\mathbf{Y}_{n}\}$ be an $\mathcal{S}_{d}^{+}$-valued supermartingale on $\{\mathcal{F}_{n}\}$, and $\tau$ be a stopping time on $\{\mathcal{F}_{n}\}$. Let $\mathbf{U}$ be a trace super-uniform random matrix taking values in $\mathcal{S}_{d}^{+}$ independent from $\mathcal{F}_{\infty}$. Then, for any $\mathbf{A}\in\mathcal{S}_{d}^{++}$,
$\mathbb{P}(\mathbf{Y}_{\tau}\npreceq\mathbf{A}^{1/2}\mathbf{U}\mathbf{A}^{1/2})\leqslant\mathrm{tr}((\mathbb{E}\mathbf{Y}_{0})\mathbf{A}^{-1}).$ (UMVI)
Consequently,
$\mathbb{P}(\exists n,\;\mathbf{Y}_{n}\npreceq\mathbf{A})\leqslant\mathrm{tr}((\mathbb{E}\mathbf{Y}_{0})\mathbf{A}^{-1}).$ (MVI)
@@
First, $\mathbb{E}\mathbf{Y}_{\tau}\preceq\mathbb{E}\mathbf{Y}_{0}$ due to Part 3 of Lemma 4.3. By Theorem 3.3 and the monotinicity of the trace,
$\mathbb{P}(\mathbf{Y}_{\tau}\npreceq\mathbf{A}^{1/2}\mathbf{U}\mathbf{A}^{1/2})\leqslant\mathrm{tr}(\mathbf{A}^{-1/2}(\mathbb{E}\mathbf{Y}_{\tau})\mathbf{A}^{-1/2})\leqslant\mathrm{tr}(\mathbf{A}^{-1/2}(\mathbb{E}\mathbf{Y}_{0})\mathbf{A}^{-1/2}),$ (27)
which is (UMVI). Next, we define the stopping time $\nu:=\inf\{n:\mathbf{Y}_{n}\npreceq\mathbf{A}\}$. Setting $\mathbf{U}$ to $\mathbf{I}$ and $\tau$ to $\nu\wedge N$ in (UMVI), we have
@@
which concludes the proof of (MVI). ∎
We remark that while there are two equivalent ways of stating the Ville’s inequality, the “any-time” statement (VI) and the stopping time statement (stated after (VI)), it appears that randomization only applies to the latter [Ramdas and Manole, 2023]. The same applies to our matrix versions of Ville’s inequality. Finally, we remark that some other maximal inequalities for scalar sub- and super-martingales can be extended to matrices as well, of which we derive a few in Appendix C.
4.2 Matrix Generalizations of Classical Test Supermartingales
Recall that scalar-valued nonnegative supermartingales are often constructed incrementally by multiplying each time a nonnegative variable with conditional expectation $\leqslant 1$. Analogously, we may use the following straightforward fact to construct $\mathcal{S}_{d}^{+}$-valued supermartingales.
Lemma 4.5.
Let $\mathbf{E}_{n}$ be $\mathcal{S}_{d}^{+}$-valued and $\mathcal{F}_{n}$-measurable, $\mathbf{F}_{n}$ be $\mathcal{S}_{d}^{++}$-valued and $\mathcal{F}_{n-1}$-measurable, such that $\mathbb{E}(\mathbf{E}_{n}|\mathcal{F}_{n-1})\preceq\mathbf{F}_{n}^{-1}$ almost surely. Then, the following $\mathcal{S}_{d}^{+}$-valued process is a supermartingale on $\{\mathcal{F}_{n}\}$:
$\mathbf{Y}_{n}=\sqrt{\mathbf{F}_{1}}\sqrt{\mathbf{E}_{1}}\sqrt{\mathbf{F}_{2}}\sqrt{\mathbf{E}_{2}}\dots\sqrt{\mathbf{F}_{n}}\sqrt{\mathbf{E}_{n}}\sqrt{\mathbf{E}_{n}}\sqrt{\mathbf{F}_{n}}\dots\sqrt{\mathbf{E}_{2}}\sqrt{\mathbf{F}_{2}}\sqrt{\mathbf{E}_{1}}\sqrt{\mathbf{F}_{1}}.$ (30)
In some cases, we can take $\mathbf{F}_{1}=\mathbf{F}_{2}=\dots=\mathbf{I}$, and thus $\mathbf{Y}_{n}=\sqrt{\mathbf{E}_{1}}\dots\sqrt{\mathbf{E}_{n}}\sqrt{\mathbf{E}_{n}}\dots\sqrt{\mathbf{E}_{1}}$.
@@ 4.2 Matrix Generalizations of Classical Test Supermartingales
We remark that by defining $\mathbf{E}_{n}^{\prime}=\sqrt{\mathbf{F}_{n}}\mathbf{E}_{n}\sqrt{\mathbf{F}_{n}}$ and $\mathbf{F}_{n}^{\prime}=\mathbf{I}$, we can obtain another $\mathcal{S}_{d}^{+}$-valued supermartingale $\sqrt{\mathbf{E}_{1}^{\prime}}\dots\sqrt{\mathbf{E}_{n}^{\prime}}\sqrt{\mathbf{E}_{n}^{\prime}}\dots\sqrt{\mathbf{E}_{1}^{\prime}}$. This option, however, involves more computation than (30) which can be computed incrementally by computing the two matrices $\sqrt{\mathbf{F}_{1}}\sqrt{\mathbf{E}_{1}}\sqrt{\mathbf{F}_{2}}\sqrt{\mathbf{E}_{2}}\dots\sqrt{\mathbf{F}_{n}}\sqrt{\mathbf{E}_{n}}$ and $\sqrt{\mathbf{E}_{n}}\sqrt{\mathbf{F}_{n}}\dots\sqrt{\mathbf{E}_{2}}\sqrt{\mathbf{F}_{2}}\sqrt{\mathbf{E}_{1}}\sqrt{\mathbf{F}_{1}}$ on the fly.
With Lemma 4.5, we can recover matrix versions of several recent “test supermartingales” (nonnegative supermartingales with initial value one) in the literature. In the following examples, we will state some conditions on a sequence of $\mathcal{S}_{d}$-valued random matrices $\{\mathbf{X}_{n}\}$ that will lead to a sequence of $\{\mathbf{E}_{n}\}$ and $\{\mathbf{F}_{n}\}$ that satisfy Lemma 4.5, so that a $\mathcal{S}_{d}^{+}$-valued supermartingale $\{\mathbf{Y}_{n}\}$ can be constructed accordingly via (30).
First, we have the MGF-type supermartingales whose well-known scalar case is mentioned by e.g. Howard et al. [2020, Fact 1(a)].
@@
Let $\{\mathbf{X}_{n}\}$ be $\mathcal{S}_{d}$-valued random matrices adapted to $\{\mathcal{F}_{n}\}$ with common conditional mean matrix $\mathbb{E}(\mathbf{X}_{n}|\mathcal{F}_{n-1})=\mathbf{M}$ and matrix-valued conditional MGF upper bound $\mathbf{G}(\gamma)\succeq\mathbb{E}(\mathrm{e}^{\gamma(\mathbf{X}_{n}-\mathbf{M})}|\mathcal{F}_{n-1})$. Then, for each $n$, we can take $\mathbf{E}_{n}=\mathrm{e}^{\gamma_{n}(\mathbf{X}_{n}-\mathbf{M})}$, $\mathbf{F}_{n}=\mathbf{G}(\gamma_{n})^{-1}$ where $\{\gamma_{n}\}$ are predictable scalars, in Lemma 4.5.
The MGF upper bounds to use may be found in Table 1. Second, the “betting” style martingales by Waudby-Smith and Ramdas [2023].
Example 4.7 (Betting).
Let $\{\mathbf{X}_{n}\}$ be $\mathcal{S}_{d}^{+}$-valued random matrices adapted to $\{\mathcal{F}_{n}\}$ with conditional means $\mathbb{E}(\mathbf{X}_{n}|\mathcal{F}_{n-1})=\mathbf{M}_{n}$ and upper bounds $\mathbf{B}_{n}$ that are predictable (i.e., $\mathbf{X}_{n}\preceq\mathbf{B}_{n}$ a.s., and $\mathbf{B}_{n}$ is $\mathcal{F}_{n-1}$-measurable). We can take, in Lemma 4.5,
@@
Besides these two examples where we extend prior scalar results to matrices, we also note that Howard et al. [2020, Section 6.5] prove a total of eight inequalities for matrix martingale differences $\{\mathbf{Z}_{n}\}$ in the style of
@@
for some respectively $\mathcal{F}_{n}$-, $\mathcal{F}_{n-1}$-measurable matrices $\mathbf{C}_{n},\mathbf{C}_{n}^{\prime}$ under eight different conditions. These directly lead to $\mathcal{S}_{d}^{+}$-valued supermartingales via Lemma 4.5. However, Howard et al. [2020] do not consider combining these conditional expectation bounds into matrix-valued supermartingales via (30). For reasons we will discuss later in Section 5; they pursue scalar supermartingales instead. Of their eight inequalities, we will introduce two interesting cases now, omitting the other six due to their similar construction.
First, the self-normalized supermartingale that only assumes the second moment exists [Howard et al., 2020, Part (f) of Section 6.5], which we (re)prove in Section B.3.
@@
Let $\{\mathbf{X}_{n}\}$ be $\mathcal{S}_{d}$-valued random matrices adapted to $\{\mathcal{F}_{n}\}$ with conditional means $\mathbb{E}(\mathbf{X}_{n}|\mathcal{F}_{n-1})=\mathbf{M}_{n}$ and conditional variance upper bounds $\mathbb{E}((\mathbf{X}_{n}-\mathbf{M}_{n})^{2}|\mathcal{F}_{n-1})\preceq\mathbf{V}_{n}$ that are predictable. We can take, in Lemma 4.5,
@@
Second, a bound for symmetric random matrices that have a symmetric (around $0\in\mathcal{S}_{d}$), potentially heavy-tailed distribution [Howard et al., 2020, Part (d) of Section 6.5] which we prove as well in Section B.3.
@@
Let $\{\mathbf{X}_{n}\}$ be $\mathcal{S}_{d}$-valued random matrices adapted to $\{\mathcal{F}_{n}\}$ with conditional means $\mathbb{E}(\mathbf{X}_{n}|\mathcal{F}_{n-1})=\mathbf{M}_{n}$ and satisfy $\mathbf{X}_{n}-\mathbf{M}_{n}|\mathcal{F}_{n-1}\stackrel{{\scriptstyle d}}{{=}}\mathbf{M}_{n}-\mathbf{X}_{n}|\mathcal{F}_{n-1}$. Then, we can take, in Lemma 4.5,
@@
Finally, let us present a novel matrix extension of the empirical Bernstein bound, in the style of Howard et al. [2021, Theorem 4], following whom we define the function
$\psi_{E}(x)=-\log(1-x)-x=\sum_{k=2}^{\infty}\frac{x^{k}}{k}$ (38)
for $x\in[0,1)$. The notation comes from the fact that $\psi_{E}$ is the log-MGF of a centered standard exponential random variable. We distinguish the following result for its novelty by labeling it a “proposition” as opposed to an “example”.
Proposition 4.10 (Empirical Bernstein Matrix Supermartingale).
@@
$\mathbf{E}_{n}=\exp(\gamma_{n}(\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n})-\psi_{E}(\gamma_{n})(\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n})^{2}),\quad\mathbf{F}_{n}=\exp(-\gamma_{n}(\mathbf{M}_{n}-\widehat{\mathbf{X}}_{n})),$ (39)

### v4

@@ 3.1 Convergence, Optional Stopping, and Maximal Inequalities
1. For any $\mathbf{v}\in\mathbb{R}^{d}$, the scalar process $\{\mathbf{v}^{\mathsf{T}}\mathbf{Y}_{n}\mathbf{v}\}$ is a supermartingale on $\{\mathcal{F}_{n}\}$ and therefore so is the stopped process $\{\mathbf{v}^{\mathsf{T}}\mathbf{Y}_{n\wedge\tau}\mathbf{v}\}$ due to the optional stopping theorem (e.g. Theorem 10.15 of Klenke [2013]). This implies that $\{\mathbf{Y}_{n\wedge\tau}\}$ is a matrix supermartingale on $\{\mathcal{F}_{n}\}$. 2. For any $\mathbf{v}\in\mathbb{R}^{d}$, we apply the optional sampling theorem (e.g. Theorem 10.11(i) of Klenke [2013]) to the scalar-valued supermartingale $\{\mathbf{v}^{\mathsf{T}}\mathbf{Y}_{n}\mathbf{v}\}$ to obtain that $\mathbf{v}^{\mathsf{T}}(\mathbb{E}\mathbf{Y}_{\tau})\mathbf{v}=\mathbb{E}(\mathbf{v}^{\mathsf{T}}\mathbf{Y}_{\tau}\mathbf{v})\leqslant\mathbb{E}(\mathbf{v}^{\mathsf{T}}\mathbf{Y}_{0}\mathbf{v})=\mathbf{v}^{\mathsf{T}}(\mathbb{E}\mathbf{Y}_{0})\mathbf{v}$. Therefore $\mathbb{E}\mathbf{Y}_{\tau}\preceq\mathbb{E}\mathbf{Y}_{0}$. 3. Applying Theorem 3.2 to the $\mathcal{S}_{d}^{+}$-valued stopped supermartingale $\{\mathbf{Y}_{\tau\wedge n}\}$, we have $\mathbb{E}\mathbf{Y}_{\tau\wedge\infty}\preceq\mathbb{E}\mathbf{Y}_{\tau\wedge 0}$, i.e., $\mathbb{E}\mathbf{Y}_{\tau}\preceq\mathbb{E}\mathbf{Y}_{0}$. ∎
This gives rise to the matrix versions of the two mutually equivalent statements of Ville’s inequality (VI) via some manipulation of crossing events.
Theorem 3.4 (Matrix Ville’s Inequality).
Let $\{\mathbf{Y}_{n}\}$ be an $\mathcal{S}_{d}^{+}$-valued supermartingale. Then, for any $\mathbf{A}\in\mathcal{S}_{d}^{++}$,
$\mathbb{P}(\exists n,\;\mathbf{Y}_{n}\npreceq\mathbf{A})\leqslant\mathrm{tr}((\mathbb{E}\mathbf{Y}_{0})\mathbf{A}^{-1});$ (MVI)
or equivalently,
for any stopping time $\text{for any stopping time }\tau,\quad\mathbb{P}(\mathbf{Y}_{\tau}\npreceq\mathbf{A})\leqslant\mathrm{tr}((\mathbb{E}\mathbf{Y}_{0})\mathbf{A}^{-1}).$ (MVI)
@@
We prove the stopped inequality first. Due to Part 3 of Theorem 3.3, $\mathbb{E}\mathbf{Y}_{\tau}\preceq\mathbb{E}\mathbf{Y}_{0}$. Now by (MMI) and the monotinicity of the trace,
$\mathbb{P}(\mathbf{Y}_{\tau}\npreceq\mathbf{A})\leqslant\mathrm{tr}(\mathbf{A}^{-1/2}(\mathbb{E}\mathbf{Y}_{\tau})\mathbf{A}^{-1/2})\leqslant\mathrm{tr}(\mathbf{A}^{-1/2}(\mathbb{E}\mathbf{Y}_{0})\mathbf{A}^{-1/2}),$ (12)
which is the stopped (MVI). Next, we define the stopping time $\nu:=\inf\{n:\mathbf{Y}_{n}\npreceq\mathbf{A}\}$. Applying the result above with $\tau=\nu\wedge N$,
@@
which concludes the proof of both forms of (MVI). ∎
A natural question arises: in comparison to the scalar (VI), does (MVI) provide tighter, looser, or incomparable concentration results? This question is essential as from every PSD supermartingale $\{\mathbf{Y}_{n}\}$ one can construct scalar nonnegative supermartingales $\{\mathrm{tr}\mathbf{Y}_{n}\}$ and $\{\mathbf{v}^{\mathsf{T}}\mathbf{Y}_{n}\mathbf{v}\}$. We shall answer the question in detail, theoretically in Sections 5.1 and 5.2 and empirically in Section 6. As our exposition unfolds, the following three conclusions shall become clear:
  1. 1.
    (MVI) is in general not reducible to the scalar (VI).
  2. 2.
    To conceptualize the difference between (MVI) and (VI), it suffices to look at the one-time case, comparing (MMI) and (MI). The spectra of both the “threshold” matrix $\mathbf{A}$ and the random matrix $\mathbf{X}$ matter.
  3. 3.
    While the one-time Markov case reveals the potential tightness benefit of (MVI), in real data the conditions for such benefit often arise naturally in the sequential, anytime-valid testing case that necessitates Ville-type inequalities.
As an aside, we present the additional result of a matrix analog of Doob’s $L^{p}$ inequality for forward matrix martingales in Section C.2. It is not necessarily relevant to our upcoming discussion in statistical contexts but might be of independent interest.
3.2 Matrix Test Supermartingales
Scalar-valued nonnegative supermartingales are often constructed incrementally: if one can find an adapted nonnegative process $\{E_{n}\}$ that has the conditional expectation bound $\mathbb{E}(E_{n}|\mathcal{F}_{n})\leqslant 1$, then their product $Y_{n}=\prod_{i=1}^{n}E_{i}$ is a nonnegative supermartingale. In particular, the supermartingale satisfies $Y_{0}=1$ and thus satisfies (VI) with $\mathbb{P}(Y_{\tau}\geqslant a)\leqslant a^{-1}$. It is called a test supermartingale. We propose the following analogous way to construct $\mathcal{S}_{d}^{+}$-valued supermartingales that involves a bit more flexibility.
Lemma 3.5 (Construction of Matrix Test Supermartingales).
If an adapted $\mathcal{S}_{d}^{+}$-valued process $\{\mathbf{E}_{n}\}_{n\geqslant 1}$ and a predictable $\mathcal{S}_{d}^{++}$-valued process $\{\mathbf{F}_{n}\}_{n\geqslant 1}$ satisfy $\mathbb{E}(\mathbf{E}_{n}|\mathcal{F}_{n-1})\preceq\mathbf{F}_{n}^{-1}$ almost surely for all $n\geqslant 1$. Then, the following is a PSD supermartingale:
$\mathbf{Y}_{n}=\sqrt{\mathbf{F}_{1}}\sqrt{\mathbf{E}_{1}}\sqrt{\mathbf{F}_{2}}\sqrt{\mathbf{E}_{2}}\dots\sqrt{\mathbf{F}_{n}}\sqrt{\mathbf{E}_{n}}\sqrt{\mathbf{E}_{n}}\sqrt{\mathbf{F}_{n}}\dots\sqrt{\mathbf{E}_{2}}\sqrt{\mathbf{F}_{2}}\sqrt{\mathbf{E}_{1}}\sqrt{\mathbf{F}_{1}},$ (15)
with $\mathbf{Y}_{0}=\mathbf{I}$. It is called a matrix test supermartingale.
In some cases, we can take $\mathbf{F}_{1}=\mathbf{F}_{2}=\dots=\mathbf{I}$, where the process has a simpler expression $\mathbf{Y}_{n}=\sqrt{\mathbf{E}_{1}}\dots\sqrt{\mathbf{E}_{n}}\sqrt{\mathbf{E}_{n}}\dots\sqrt{\mathbf{E}_{1}}$.
@@ 3.2 Matrix Test Supermartingales
The expression of $\mathbf{Y}_{n}$ involves splitting the stepwise matrices into square roots and multiplying bidirectionally from the outside, which, we note, ensures the product is still in $\mathcal{S}_{d}$ and has the conditional expectation property we want; in particular, $\mathbf{F}_{n}\mathbf{E}_{n}$ and $\mathbf{E}_{n}\mathbf{F}_{n}$ may not be symmetric. The process $\{\mathbf{Y}_{n}\}$ can be computed incrementally on the fly by noting that
$\mathbf{Y}_{n}=\mathbf{L}_{n}\mathbf{R}_{n},\quad\mathbf{L}_{n}=\mathbf{L}_{n-1}\sqrt{\mathbf{F}_{n}}\sqrt{\mathbf{E}_{n}},\quad\mathbf{R}_{n}=\sqrt{\mathbf{E}_{n}}\sqrt{\mathbf{F}_{n}}\mathbf{R}_{n-1}.$ (18)
We remark that by defining $\mathbf{E}_{n}^{\prime}=\sqrt{\mathbf{F}_{n}}\mathbf{E}_{n}\sqrt{\mathbf{F}_{n}}$ and $\mathbf{F}_{n}^{\prime}=\mathbf{I}$, we can obtain another PSD supermartingale $\sqrt{\mathbf{E}_{1}^{\prime}}\dots\sqrt{\mathbf{E}_{n}^{\prime}}\sqrt{\mathbf{E}_{n}^{\prime}}\dots\sqrt{\mathbf{E}_{1}^{\prime}}$. This option, however, involves slightly more matrix product computation than (15).
Algorithm 1 Sequential Test via a Matrix Test Supermartingale
Null hypothesis $\mathcal{H}_{0}$ on the data stream $X_{1},X_{2},\dots$ taking values in $\mathcal{X}$;
Functions $\mathbf{E}:\mathcal{X}\to\mathcal{S}_{d}^{+}$ and $\mathbf{F}:\mathcal{X}\to\mathcal{S}_{d}^{++}$ such that $\mathbb{E}(\mathbf{E}(X_{n})|X_{1},\dots,X_{n-1})\preceq\mathbf{F}^{-1}(X_{n})$ whenever $\mathcal{H}_{0}$ is true;
Type-I error level $\alpha\in(0,1)$; $\mathbf{A}\in\mathcal{S}_{d}^{++}$ such that $\mathrm{tr}(\mathbf{A}^{-1})=\alpha$
$\mathcal{H}_{0}$ is falsely rejected with probability at most $\alpha$
$\mathbf{L}\leftarrow\mathbf{I}$, $\mathbf{R}\leftarrow\mathbf{I}$
for $n=1,2,\dots$ do
     observe $X_{n}$
     $\mathbf{L}\leftarrow\mathbf{L}\sqrt{\mathbf{F}(X_{n})}\sqrt{\mathbf{E}(X_{n})}$
     $\mathbf{R}\leftarrow\sqrt{\mathbf{E}(X_{n})}\sqrt{\mathbf{F}(X_{n})}\mathbf{R}$
     if  $\mathbf{L}\mathbf{R}\npreceq\mathbf{A}$  then
         reject $\mathcal{H}_{0}$
     end if
end for
With Lemma 3.5, we can recover matrix versions of several scalar test supermartingales in the recent sequential statistics literature. In the next subsection, we will state some conditions on a sequence of $\mathcal{S}_{d}$-valued random matrices $\{\mathbf{X}_{n}\}$ that will lead to two sequences of $\{\mathbf{E}_{n}\}$ and $\{\mathbf{F}_{n}\}$ that satisfy Lemma 3.5, so that a $\mathcal{S}_{d}^{+}$-valued supermartingale $\{\mathbf{Y}_{n}\}$ can be constructed accordingly via (15). Combining with (MVI), we form the following generic scheme for sequentially testing a null hypothesis $\mathcal{H}_{0}$ using a PSD supermartingale outlined in Algorithm 1, where if $\mathcal{H}_{0}$ is true the procedure has at most $\alpha$ probability ever rejecting it.
We further remark that, while with scalar nonnegative supermartingales formed multiplicatively, one often works on the log scale to ease the computation. That is, if $M_{n}=\prod_{i=1}^{n}E_{i}$ is a scalar nonnegative supermartingale, one often adds up $\log E_{1}+\dots+\log E_{n}$ and rejects the null when this sum exceeds $\log(1/\alpha)$. With the PSD supermartingale formed in Lemma 3.5, while it is indeed true that
[…]
due to the monotonicity of the matrix logarithm (Fact 1), one can not compute $\log\mathbf{Y}_{n}$ via summing $\log\mathbf{E}_{i}$ and $\log\mathbf{F}_{i}$. An additive alternative to (MVI) that avoids matrix multiplication might be of interest for future work.
3.3 Examples: Conditional Mean Testing Problems
We have numerous examples for Lemma 3.5 where one can conduct sequential hypothesis testing with Algorithm 1. All examples below, we note, involve the null hypothesis in the form of “the conditional mean of $\mathbf{X}_{n}$ equals $\mathbf{M}_{n}$” where $\{\mathbf{M}_{n}\}$ is a pre-specified predictable sequence. A common special case is when the matrices $\{\mathbf{X}_{n}\}$ are i.i.d. and one tests if their common mean equals a specific value.
First, we have the moment generating function (MGF)-based supermartingales whose well-known scalar case is mentioned by e.g. Howard et al. [2020, Fact 1(a)].
@@ 3.3 Examples: Conditional Mean Testing Problems
Let $\{\mathbf{X}_{n}\}$ be an adapted sequence of $\mathcal{S}_{d}$-valued random matrices with conditional means $\mathbb{E}(\mathbf{X}_{n}|\mathcal{F}_{n-1})=\mathbf{M}_{n}$ and matrix-valued conditional MGF upper bound $\mathbf{G}_{n}(\gamma)\succeq\mathbb{E}(\mathrm{e}^{\gamma(\mathbf{X}_{n}-\mathbf{M}_{n})}|\mathcal{F}_{n-1})$. Then, for each $n$, we can take $\mathbf{E}_{n}=\mathrm{e}^{\gamma_{n}(\mathbf{X}_{n}-\mathbf{M}_{n})}$, $\mathbf{F}_{n}=\mathbf{G}_{n}(\gamma_{n})^{-1}$ where $\{\gamma_{n}\}$ are predictable scalars, in Lemma 3.5.
It is worth noting that in general it is hard to prove matrix MGF bounds and they are known only in very few cases. On the other hand, much more is known on bounding the matrix log-MGF $\log\left(\mathbf{G}_{n}(\cdot)\right)$ under various distribution assumptions, obtained via the Lieb-Tropp technique mentioned in Section 2.3.1. These log-MGF upper bounds only lead to scalar-valued supermartingales (9) instead of matrix-valued supermartingales. We review some existing results of matrix MGFs by Tropp [2012, 2015] in Section C.1.
Our second example extends the wealth process of betting against the fairness of a coin. The scalar case was initially proposed for Bernoulli coin tosses, generalized to any bounded random scalars by Shafer and Vovk [2005, Chapters 3.1 and 3.2 resp.], and recently applied to sequential mean testing by Waudby-Smith and Ramdas [2023]. We have the following corresponding result for bounded random matrices.
Example 3.7 (Matrix Betting).
Let $\{\mathbf{X}_{n}\}$ be an adapted sequence of $\mathcal{S}_{d}^{+}$-valued random matrices with conditional means $\mathbb{E}(\mathbf{X}_{n}|\mathcal{F}_{n-1})=\mathbf{M}_{n}$ and upper bounds $\mathbf{B}_{n}$ that are predictable (i.e., $\mathbf{X}_{n}\preceq\mathbf{B}_{n}$ a.s., and $\mathbf{B}_{n}$ is $\mathcal{F}_{n-1}$-measurable). We can take, in Lemma 3.5,
@@
Besides these two examples where we extend prior scalar results to matrices, we also note that Howard et al. [2020, Section 6.5] prove a total of eight inequalities for $\mathcal{S}_{d}$-valued martingale differences $\{\mathbf{Z}_{n}\}$ in the style of
@@
for some respectively $\mathcal{F}_{n}$-, $\mathcal{F}_{n-1}$-measurable matrices $\mathbf{C}_{n},\mathbf{C}_{n}^{\prime}$ under eight different conditions. These directly lead to $\mathcal{S}_{d}^{+}$-valued supermartingales via Lemma 3.5. However, Howard et al. [2020] do not consider combining these conditional expectation bounds into matrix-valued supermartingales via (15); they pursue scalar supermartingales instead, with methods we discuss later in Section 5.2. Of their eight inequalities, we will introduce two interesting cases now and (re-)prove them in Section B.2, omitting the other six due to their similar construction.
First, the self-normalized supermartingale that only assumes the second moment exists [Howard et al., 2020, Part (f) of Section 6.5].
@@
Let $\{\mathbf{X}_{n}\}$ be an adapted sequence of $\mathcal{S}_{d}$-valued random matrices with conditional means $\mathbb{E}(\mathbf{X}_{n}|\mathcal{F}_{n-1})=\mathbf{M}_{n}$ and conditional variance upper bounds $\mathbb{E}((\mathbf{X}_{n}-\mathbf{M}_{n})^{2}|\mathcal{F}_{n-1})\preceq\mathbf{V}_{n}$ that are predictable. We can take, in Lemma 3.5,
@@
Second, a bound for symmetric random matrices that have a symmetric (around $0\in\mathcal{S}_{d}$), potentially heavy-tailed distribution [Howard et al., 2020, Part (d) of Section 6.5].
@@
Let $\{\mathbf{X}_{n}\}$ be an adapted sequence of $\mathcal{S}_{d}$-valued random matrices with conditional means $\mathbb{E}(\mathbf{X}_{n}|\mathcal{F}_{n-1})=\mathbf{M}_{n}$ and satisfy $\mathbf{X}_{n}-\mathbf{M}_{n}|\mathcal{F}_{n-1}\stackrel{{\scriptstyle d}}{{=}}\mathbf{M}_{n}-\mathbf{X}_{n}|\mathcal{F}_{n-1}$. Then, we can take, in Lemma 3.5,
@@
Further, let us present two examples in the light- and heavy-tail regimes respectively that do not appear among the ones due to Howard et al. [2020, Section 6.5]. First, a matrix extension of empirical Bernstein inequalities, in the style of Howard et al. [2021, Theorem 4], following whom we define the function
$\psi_{\mathrm{E}}(x)=-\log(1-x)-x$ (25)
for $x\in[0,1)$. The notation comes from the fact that $\psi_{\mathrm{E}}$ is the log-MGF of a centered standard exponential random variable. The following is a rephrasing of a recent result by Wang and Ramdas [2024, Lemma 4.1] in the language of our Lemma 3.5.
Example 3.10 (Empirical Bernstein-Type Matrix Supermartingale).
@@
$\mathbf{E}_{n}=\exp(\gamma_{n}(\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n})-\psi_{\mathrm{E}}(\gamma_{n})(\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n})^{2}),\quad\mathbf{F}_{n}=\exp(-\gamma_{n}(\mathbf{M}_{n}-\widehat{\mathbf{X}}_{n})),$ (26)